# Unsupervised Detection of Spatiotemporal Anomalies in PMU Data Using Transformer-Based BiGAN

### v1

@@ I-A Background and Literature Review
The electric power grid is rapidly evolving into a complex cyber-physical ecosystem, driven by the proliferation of distributed energy resources, advanced sensing devices, and real-time communication networks. Although these innovations promise enhanced flexibility and efficiency, they also introduce novel security and reliability challenges. In particular, phasor measurement units (PMUs) - which provide high-resolution, time-synchronized voltage and current phasors - have become indispensable for situational awareness, control, and protection in modern grids [1].
@@ I-C Transformer-Based and Unsupervised Architectures
Beyond GANs, attention-driven and transformer architectures [7] excel at modeling long-range temporal dependencies and spatial correlations inherent in wide-area monitoring data. Self-attention layers allow the network to dynamically weight phasor channels, allowing focus on key features across temporal and spatial dimensions [3]. The successful detection of anomalies in high-resolution PMU streams has also been achieved using other approaches [8]. These include methods such as autoencoders with LSTM encoders [9] and density-based clustering algorithms [10].
@@ I-D Realistic PMU Dataset and Benchmarking
We used the publicly accessible labeled PMU dataset generated on a hardware-in-the-loop (HIL) cyber–power testbed, as detailed by Mustafa et al. [1]. This testbed integrates real-time digital simulators, hardware, and software phasor measurement units (PMUs), remote terminal units, a cloud-based aggregation service, and a network emulator to recreate operational grid conditions and inject controlled anomaly events.
@@ III-A Data Pre-processing
Compared with Min–Max, global log, or Robust/Quantile scalers (which either preserve skew or flatten informative tails), this selective-log + standardization preserves angles/frequency semantics, equalizes scales without leakage, and yielded more stable training and better early PR/ROC in our setting(Fig. 2).
@@ III-F Auxiliary Loss Terms
All numeric PMU channels (F=112F{=}112) are reconstructed jointly for each $T$-step window $\mathbf{x}\!\in\!\mathbb{R}^{T\times F}$, and losses average over time and features. The reconstruction loss used in training is mean absolute error (MAE):
@@ III-G Anomaly Scoring
For evaluation, reconstruction residuals are feature–weighted to reflect channel variability. Let $w_{f}\propto\operatorname{Var}_{\text{train}}(x_{:,f})^{-1}$ computed on normal training data and normalized so $\frac{1}{F}\sum_{f}\tilde{w}_{f}=1$. The window score is
@@ III-I Training Details
The proposed Transformer-augmented BiGAN (T-BiGAN) is optimized with Adam [14] (lr $3\times 10^{-4}$, $\beta_{1}=0.5$, $\beta_{2}=0.999$), batch size 64, and alternating updates of the encoder–generator and discriminator; latent vectors $\mathbf{z}\!\sim\!\mathcal{N}(\mathbf{0},\mathbf{I})$. Hyperparameters are chosen via a small random search on a held-out validation split (5 trials/setting, selecting by mAP with ROC–AUC as tie-breaker) over: lr $1\!\times\!10^{-5},5\!\times\!10^{-4}]$, dropout $\{0,0.1,0.2\}$, $\lambda_{\text{rec}}\!\in\!\{1,5,10,25\}$, $\lambda_{z}\!\in\!\{0,0.5,1,2\}$, $\alpha\!\in\!\{0.4,0.6,0.8\}$, label smoothing $\{1.0,0.9\}$, spectral norm $\{\text{off},\text{on}\}$, and gradient penalty $\{0,10\}$. Spectral normalization [15] and mild label smoothing ($0.9$) stabilized discriminator training; larger $\lambda_{\text{rec}}$ increased precision at some recall cost, while dropout >0.2>0.2 degraded recall; with spectral norm enabled, gradient penalty was typically unnecessary. We train for 50 epochs on a single NVIDIA Quadro RTX 8000 GPU, with convergence typically by epochs 35–40.
@@ IV-A Evaluation Protocol
We benchmark T-BiGAN on the realistic, labeled PMU dataset from Hussain et al.[16]. Frame‑level anomaly detection performance is measured using Precision, Recall, F1‑score, and ROC‑AUC. Thresholds are chosen via Youden’s JJ‑statistic [17] on the validation split. Baselines include PCA [18], Isolation Forest [19], Autoencoder (AE), Variational Autoencoder (VAE), LSTM‑AE [20], BiLSTM-AE [21] and a Transformer + GAN  [22] approach.
@@ IV-B Quantitative Comparison
TABLE I: Detection performance (mean ± ς over 5 seeds) on the Hussain et al. PMU dataset. The best figures per column are bold.
@@
PCA - $0.58$ p m 0.02 $0.71$ p m 0.03 $0.64$ p m 0.02 $0.79$ p m 0.01
Isolation Forest - $0.61$ p m 0.02 $0.75$ p m 0.03 $0.67$ p m 0.04 $0.82$ p m 0.02
AE $0.84$ $0.70$ p m 0.03 $0.63$ p m 0.04 $0.66$ p m 0.05 $0.85$ p m 0.03
VAE $0.86$ $0.72$ p m 0.04 $0.66$ p m 0.05 $0.69$ p m 0.09 $0.86$ p m 0.07
LSTM–AE $0.88$ $0.75$ p m 0.07 $0.70$ p m 0.08 $0.72$ p m 0.03 $0.88$ p m 0.05
BiLSTM–AE $0.23$ $0.82$ p m 0.07 $0.76$ p m 0.08 $0.80$ p m 0.03 $0.89$ p m 0.06
Transformer + GAN $1.46$ $0.88$ p m 0.12 $0.77$ p m 0.14 $0.82$ p m 0.13 $0.91$ p m 0.11
BiGAN + CNN $1.43$ $0.87$ p m 0.06 $0.78$ p m 0.07 $0.85$ p m 0.06 $0.90$ p m 0.12
Transformer + BiGAN (no attn) $1.41$ $0.88$ p m 0.14 $0.79$ p m 0.12 $0.83$ p m 0.21 $0.92$ p m 0.09
Transformer–attn–BiGAN (Ours) $0.99$ $0.90\pm 0.04$ $1\pm 0.03 0.88±0$$8\pm 0.05$ $0.95\pm 0.02$
  • “-” denotes detectors that are intrinsically threshold‐free; we operate them at the score quantile that maximizes Youden’s $J$ on the validation split. All other thresholds are fixed on validation and not re-optimized on test data.
Table I summarizes mean ± ς (over 5 seeds) for each metric. Our T‑BiGAN achieves a precision of 0.90 ± 0.04, recall of 0.81 ± 0.03, F1‑score of 0.88 ± 0.05, and ROC‑AUC of 0.95 ± 0.02, outperforming all the other baselines.
@@ IV-D Confusion Matrix Analysis
To further characterize real-time performance, we visualize the confusion matrix obtained at the operating threshold $\theta=0.9902$ (Fig. 6). The matrix is computed over the test set of 44,000 frames.
From Fig. 6:
  • True Negatives (TN): 17 481
  • False Positives (FP): 1 422
  • False Negatives (FN): 446
  • True Positives (TP): 11
These counts correspond to:
$\mathrm{TPR}=\frac{11}{11+446}=0.024,\quad\mathrm{FPR}=\frac{1\,422}{1\,422+17\,481}=0.075.$
Even with high precision, the low prevalence of anomalies leads to a small true-positive rate on a per-frame basis, highlighting the challenge of detecting rare events in streaming PMU data.
Figure 6: Confusion matrix of T-BiGAN on the test set at threshold $\theta=0.9902$.
@@ IV-E Loss Convergence and Anomaly‑Score Dynamics
Figure 7: (Up) Generator, discriminator, and encoder loss curves. (Down) Corresponding anomaly-score trace over time.
@@ VI Conclusion
We have presented T-BiGAN, a Transformer-augmented BiGAN that consistently outperforms traditional and deep-learning-based detectors on a high-fidelity PMU dataset. Our approach excels at uncovering both abrupt frequency excursions and faint voltage oscillations, thanks to the global attention mechanism’s ability to model long-range dependencies. By aligning data and latent spaces through a joint discriminator and cycle-consistency loss, T-BiGAN achieves an ROC-AUC of $\mathbf{0.9956}$ and an F1-score of $\mathbf{0.8862}$, while its adaptive thresholding keeps false alarms at bay even as system conditions evolve. These results demonstrate that combining self-attention with bidirectional adversarial learning not only elevates detection accuracy but also enhances robustness, key requirements for real-time grid monitoring and edge-level deployment.

### v2

@@ I-A Background and Literature Review
The electric power grid is rapidly evolving into a complex cyber-physical ecosystem, driven by the proliferation of distributed energy resources, advanced sensing devices, and real-time communication networks. Although these innovations promise enhanced flexibility and efficiency, they also introduce novel security and reliability challenges. Phasor measurement units (PMUs) offer high-resolution, time-synchronized voltage and current phasors, making them essential for situational awareness, control, and protection in contemporary power grids [18].
@@ I-C Transformer-Based and Unsupervised Architectures
Beyond GANs, attention-driven and transformer architectures [22] excel at modeling long-range temporal dependencies and spatial correlations inherent in wide-area monitoring data. Self-attention layers allow the network to dynamically assign weights to phasor channels, facilitating the emphasis on critical features in both temporal and spatial dimensions [5]. Successful detection of anomalies in high-resolution PMU streams has also been achieved using other approaches [3]. These include methods such as autoencoders with LSTM encoders [7] and density-based clustering algorithms [14].
@@ I-D Realistic PMU Dataset and Benchmarking
We used the publicly accessible labeled PMU dataset generated on a hardware-in-the-loop (HIL) cyber-power testbed, as detailed by Mustafa et al. [18]. This testbed integrates real-time digital simulators, hardware and software phasor measurement units (PMUs), remote terminal units, a cloud-based aggregation service, and a network emulator to recreate operational grid conditions and inject controlled anomaly events.
@@ III-A Data Pre-processing
Compared with Min–Max, global log, or Robust/Quantile scalers (which either preserve skew or flatten informative tails), this selective-log + standardization preserves angles/frequency semantics, equalizes scales without leakage, and yields more stable training and better early PR/ROC in our setting (Fig. 2).
@@ III-F Auxiliary Loss Terms
All numeric PMU channels (F=112F{=}112) are reconstructed jointly for each $T$-step window $\mathbf{x}\!\in\!\mathbb{R}^{T\times F}$, and the losses are averaged over time and features. The reconstruction loss used in training is mean absolute error (MAE):
@@ III-G Anomaly Scoring
For evaluation, reconstruction residuals are feature-weighted to reflect channel variability. Let, $w_{f}\propto\operatorname{Var}_{\text{train}}(x_{:,f})^{-1}$ computed on normal training data and normalized so $\frac{1}{F}\sum_{f}\tilde{w}_{f}=1$. The window score is
@@ III-I Training Details
The proposed Transformer-augmented BiGAN (T-BiGAN) is optimized with Adam [12] (lr $3\times 10^{-4}$, $\beta_{1}=0.5$, $\beta_{2}=0.999$), batch size 64, and alternating updates of the encoder–generator and discriminator; latent vectors $\mathbf{z}\!\sim\!\mathcal{N}(\mathbf{0},\mathbf{I})$. Hyperparameters are chosen via a small random search on a held-out validation split (5 trials/setting, selecting by mAP with ROC–AUC as a tie-breaker) over: lr $1\!\times\!10^{-5},5\!\times\!10^{-4}]$, dropout $\{0,0.1,0.2\}$, $\lambda_{\text{rec}}\!\in\!\{1,5,10,25\}$, $\lambda_{z}\!\in\!\{0,0.5,1,2\}$, $\alpha\!\in\!\{0.4,0.6,0.8\}$, label smoothing $\{1.0,0.9\}$, spectral norm $\{\text{off},\text{on}\}$, and gradient penalty $\{0,10\}$. Spectral normalization [17] and mild label smoothing ($0.9$) stabilized discriminator training; the larger $\lambda_{\text{rec}}$ increased precision at some recall cost, while dropout >0.2>0.2 reduced recall; with the spectral norm enabled, the gradient penalty was typically unnecessary. We train for 50 epochs on a single NVIDIA Quadro RTX 8000 GPU, with convergence typically by epochs 35–40.
@@ IV-A Evaluation Protocol
We benchmark T-BiGAN on the realistic, labeled PMU dataset from Hussain et al.[11]. Frame‑level anomaly detection performance is measured using Precision, Recall, F1‑score, and ROC‑AUC. Thresholds are chosen via Youden’s JJ‑statistic [25] on the validation split. Baselines include PCA [1], Isolation Forest [13], Autoencoder (AE), Variational Autoencoder (VAE), LSTM‑AE [16], BiLSTM-AE [19], and a Transformer + GAN  [24] approach.
@@ IV-B Quantitative Comparison
TABLE I: Detection performance (mean ± ς over 5 seeds) on the Hussain et al. [11] PMU dataset. The best figures per column are bold.
@@
PCA $0.58\pm 0.02$ $0.71\pm 0.03$ $0.64\pm 0.02$ $0.79\pm 0.01$
Isolation Forest $0.61\pm 0.02$ $0.75\pm 0.03$ $0.67\pm 0.04$ $0.82\pm 0.02$
AE 0.84 $0.70\pm 0.03$ $3\pm 0.04 0.66±0$$6\pm 0.05$ $0.85\pm 0.03$
VAE 0.86 $0.72\pm 0.04$ $0.66\pm 0.05$ $0.69\pm 0.09$ $0.86\pm 0.07$
LSTM–AE 0.88 $0.75\pm 0.07$ $0.70\pm 0.08$ $0.72\pm 0.03$ $0.88\pm 0.05$
BiLSTM–AE 0.23 $0.82\pm 0.07$ $0.76\pm 0.08$ $0.80\pm 0.03$ $0.89\pm 0.06$
Transformer + GAN 0.89 $0.88\pm 0.12$ $0.77\pm 0.14$ $0.82\pm 0.13$ $0.91\pm 0.11$
BiGAN + CNN 0.86 $0.87\pm 0.06$ $0.78\pm 0.07$ $0.85\pm 0.06$ $0.90\pm 0.12$
Transformer + BiGAN (no attn) 0.88 $0.88\pm 0.14$ $0.79\pm 0.12$ $0.83\pm 0.21$ $0.92\pm 0.09$
Transformer–attn–BiGAN (Ours) 0.93 $0.90\pm 0.04$ $1\pm 0.03 0.88±0$$8\pm 0.05$ $0.95\pm 0.02$
  • “–” denotes detectors that are intrinsically threshold-free; we operate them at the score quantile that maximizes Youden’s $J$ on the validation split. All other thresholds are fixed on validation and not re-optimized on test data.
Table I summarizes mean ± ς (over 5 seeds) for each metric. Our T‑BiGAN achieves a precision of 0.90 ± 0.04, a recall of 0.81 ± 0.03, an F1‑score of 0.88 ± 0.05, and ROC‑AUC of 0.95 ± 0.02, outperforming all the other baselines.
@@ IV-D Confusion Matrix Analysis
Fig. 6 shows the confusion matrix on the test set (44,000 frames) at the optimized threshold $\theta=0.9902$. The model achieves a strong balance between detection and false-alarm control:
  • TN: 17 472, FP: 9
  • FN: 19, TP: 81
$\mathrm{Precision}=0.90,\quad\mathrm{Recall}=0.81.$
The results indicate robust detection of rare events with minimal false positives, confirming reliable real-time performance in PMU streams.
Figure 6: Confusion matrix of T-BiGAN at $\theta=0.9902$ (TP highlighted in light blue).
@@ IV-E Loss Convergence and Anomaly‑Score Dynamics
Figure 7: (Up) Generator, discriminator, and encoder loss curves for our model. (Down) Anomaly-score trace over time for different models.
@@ VI Conclusion
We have presented T-BiGAN, a Transformer-augmented BiGAN that consistently outperforms traditional and deep-learning-based detectors on a high-fidelity PMU dataset. Our approach excels at uncovering both abrupt frequency excursions and faint voltage oscillations, thanks to the global attention mechanism’s ability to model long-range dependencies. By aligning data and latent spaces through a joint discriminator and cycle-consistency loss, T-BiGAN achieves an ROC-AUC of $\mathbf{0.9956}$ and an F1-score of $\mathbf{0.8862}$, while its adaptive thresholding keeps false alarms at bay even as system conditions evolve. These results demonstrate that combining self-attention with bidirectional adversarial learning not only elevates detection accuracy but also enhances robustness, key requirements for real-time grid monitoring and edge-level deployment. Our results are specific to the dataset  [11] and may require further study to confirm their applicability to other datasets.